Find three consecutive even integers such that twice the sum of the first and third is four more than three times the second. Thank You
Answer by mananth(16946)

let the numbers be x, x+2, x+4
2(x+x+4)= 3(x+2)+4
2(2x+4)=3x+6+4
4x+8 = 3x +10
add -8 to both sides
4x+8-8=3x+10-8
4x=3x+2
add -3x to both sides
4x-3x=3x-3x+2
x=2
the numbers are 2,4,6